OpenStudy (anonymous):

Alvin paddled for 3 hours with a 5-km/h current to reach a campsite. The return trip against the same current took 8 hours. Find the speed of the boat in still water.

OpenStudy (anonymous):

d1 = 3 ( v + 5 ) d2 = 8 ( v -5) Thus: 3 ( v + 5 ) = 8 ( v -5) @katiep24 Can you solve it from here?

OpenStudy (anonymous):

Yes, thank you so much